1) A 10-year corporate bond has a coupon rate of 6% with annual payments. If the current value of the bond in the marketplace is $900, then what is the Yield-to-Maturity (YTM)?

2) A 10-year corporate bond has a coupon rate of 6% with annual payments. If the current value of the bond in the marketplace is $1100, then what is the Yield-to-Maturity (YTM)?

3) A 10-year corporate bond has a coupon rate of 6% with semi-annual payments. If the current value of the bond in the marketplace is $900, then what is the Yield-to-Maturity (YTM)?
Answer #1

1)

Yield to maturity of a bond is the internal rate of return earned by investor.

Bond price is the initial year cash outflow.

Periodic cash flow = Face value x coupon rate = $1,000 x 0.06 = $ 60

Last period cash flow = Face value + coupon amount = $ 1,000 + $ 60 = $ 1,060

IRR can be computed using excel as:

A

B

1

Year

Cash flow

2

0

$          (900)

3

1

$                  60

4

2

$                  60

5

3

$                  60

6

4

$                  60

7

5

$                  60

8

6

$                  60

9

7

$                  60

10

8

$                  60

11

9

$                  60

12

10

$             1,060

13

IRR

7.45%

If above table is excel sheet, use formula “=IRR(B2:B12) in cell B13 to get IRR as 14.96%

Yield -to-Maturity of the bond is 7.45 %

2)

Yield to maturity of a bond is the internal rate of return earned by investor.

Bond price is the initial year cash outflow.

Periodic cash flow = Face value x coupon rate = $1,000 x 0.06 = $ 60

Last period cash flow = Face value + coupon amount = $ 1,000 + $ 60 = $ 1,060

IRR can be computed using excel as:

A

B

1

Year

Cash flow

2

0

$ (1,100)

3

1

$                  60

4

2

$                  60

5

3

$                  60

6

4

$                  60

7

5

$                  60

8

6

$                  60

9

7

$                  60

10

8

$                  60

11

9

$                  60

12

10

$             1,060

13

IRR

4.72%

If above table is excel sheet, use formula “=IRR(B2:B12) in cell B13 to get IRR as 4.72%

Yield -to-Maturity of the bond is 4.72 %

3)

Yield to maturity of a bond is the internal rate of return earned by investor.

Bond price is the initial year cash outflow.

Periodic cash flow = Face value x coupon rate = $1,000 x 0.06/2 = $ 1,000 x 0.03 = $ 30

Last period cash flow = Face value + coupon amount = $ 1,000 + $ 30 = $ 1,030

IRR can be computed using excel as:

A

B

1

Period

Cashflow

2

0

$ (900)

3

1

$         30

4

2

$         30

5

3

$         30

6

4

$         30

7

5

$         30

8

6

$         30

9

7

$         30

10

8

$         30

11

9

$         30

12

10

$         30

13

11

$         30

14

12

$         30

15

13

$         30

16

14

$         30

17

15

$         30

18

16

$         30

19

17

$         30

20

18

$         30

21

19

$         30

22

20

$   1,030

23

IRR

3.72%

If above table is excel sheet use formula “=IRR(B2:B22) in cell B23 to get IRR as 3.72%

Yield -to-Maturity of the bond = 3.72 % x 2 = 7.44 %